# Combining Tests and Proofs for Better Software Verification

### v1

@@ Abstract.
copyright: acmlicensedjournalyear: 2025isbn: 978-1-4503-XXXX-X/2018/06ccs: Software and its engineering Error handling and recoveryccs: Software and its engineering Software verification and validationccs: Software and its engineering Software testing and debugging
@@ 1. From opposition to complementarity
  • ‘Seeding Contradiction” (Section 6 and (Huang et al., 2023)), extends the capability to systematic test generation. By deliberately creating failed proofs and running them through Proof2Test, it can produce a collection of tests that satisfy code coverage criteria. It also incorporates loop unrolling to create tests that explore loop behavior and helped provide an assessment of MC/DC, a testing criterion sometimes viewed skeptically but widely used in industry.
@@ 2. The core idea: proofs and their failures
A prover works on the combination of a program and a specification (often expressed as “contracts” as explained in the next section). The starting point for the present work is to take advantage of the technique used internally by many modern provers: they seek one or more “counterexamples” — example inputs falsifying the expected properties If this attempts fails, the proof succeeds [^1] The duality between failure and success is a general theme of the present work. It follows from the generalization of De Morgan’s laws to predicate calculus: the negation of a universal is the existential of the negation, and conversely. In symbols, $\neg(\forall x\mid p)=\exists x\mid\neg p$. A proof seeks a universal ($\forall$) property; a failed test is an existential ($\exists$) property of the negation.. The present work focuses on the intermediate steps, when the proof fails, meaning the search for a counterexample succeeds; and, more specifically on the counterexamples themselves.
@@
The next application is to use the counterexamples to generate program fixes. Automatic program repair (APR) needs “invariants” characterizing failed executions. Instead of getting them from manually generated tests, the present approach can generate guaranteed invariants; instead of validating fixes through tests, it can guarantee their correctness by running them through the prover. Those are the functions of Proof2Fix (Section 5)
@@ 4. Creating tests from failed proofs
In Fig.  3 (a) the values are small: a two-element array with values 0 and 1. If asked to produce one counterexample, the SMT solver typically yields much larger values. Here it initially produces an array size a.count of 11,800 and values a [11799] ==\, 0 and a [11800] ==\, 5. For disproof purposes (showing that the postcondition does not hold), they are as good as any others. Since the purpose of Proof2Test is to provide concrete evidence helping programmers figure out concretely what is going on, we prefer values that mean something at the human scale. Proof2Test takes advantage of the underlying prover’s ability to generate not just one but several counterexamples (when provided with different seeds). It applies a minimization algorithm (Huang et al., 2022) to generate successive ones until it finds minimal values, such as those of Fig.  3. As with traditional programmer-devised test cases, they enable the programmer to relate to the failure by trying them out, as in Fig. 2, and to see that on loop initialization at lines 10 – 12 Result = 0 and i = 2 so that at line 13 the exit condition evaluates to True, terminating the loop and revealing the fault: early termination prevents the program from getting to the maximum value at position 2. To eliminate the error, it suffices to strengthen the exit condition to permit one more loop iteration: change i $\geq$ a.count to i >> a.count.
@@
Experiment results (Huang et al., 2022) also show that minimization is cost-effective: in most cases, it reduces the values of integer variables by over 80% with an average cost of less than 4 extra reverification runs. Most minimized values are small and relatable: out of 125, 108 are in the range [-2, 2], out of which 58 are zero; others are usually close to some constants appearing the program.
@@ 7. Related work
The earliest work we know to have applied this idea (Angeletti et al., 2009) generates tests for low-level C programs using Bounded Model Checking (BMC) (Kroening and Tautschnig, 2014), producing test suites with exhaustive branch coverage. A more recent variant, for Java bytecode, is JBMC (Brenguier et al., 2023), where each verification run only activates one assertion at a time, producing one counterexample (the C version (Angeletti et al., 2009) uses compile-time macros, one for each block to avoid actually generating multiple programs); in contrast, the SC framework produces only one program, using a single run-time variable representing the block number. Similarly, (Beyer et al., 2004) extends the software model checker Blast to generate test suites of full coverage with respect to given predicates. BMC-based approaches rely on the correctness of the bound of the execution trace: if the bound is not set correctly, some branches might not be covered, requiring more verification runs to obtain a better bound. The SC approach does not need such mechanisms. An approach (Nilizadeh et al., 2022) that applies ideas of fault injection for generating tests (building on work on using counterexamples for program repair (Nilizadeh et al., 2021)) in Hoare-style verification exploits counterexamples produced by the OpenJML (Cok, 2021) verification tool to generate unit tests in JUnit format. These approaches all bear similarities to the present work, which seems to provide the most far-ranging and unifying collection of proof-test combinations techniques, from exploiting proof failures to counterexample minimization, test generation covering several code coverage measures and automatic program repair.
Automated test generation is an active research area (e.g. (Gorla et al., 2025)) which in recent years has increasingly used LLM, generally leading to average coverage levels of between 75% to 90%. Proof2Test is much more efficient. LLMs can also serve to generate specifications automatically (Ma et al., 2025), although questions remain: while such specifications might make the code provable, they might also be the wrong specifications, preventing the generation of relevant some tests.
The present work distinguishes itself by the use of Eiffel. While it can be viewed as a limitation, this choice also provides a considerable advantage since it does require an addition to another language (as JML for Java) but uses an existing (ISO-standardized) language, taking advantage of the built-in contract mechanism and hence directly using the original contract-equipped source code for both tests (with EiffelStudio’s mechanism for evaluating assertions during testing) and proofs. Users of other language may view the present results as those of “laboratory work”, using ideal conditions so as to open the way for others to extend the results to other environments.

### v2

@@ Abstract.
copyright: acmlicensedjournalyear: 2026isbn: 978-1-4503-XXXX-X/2018/06ccs: Software and its engineering Error handling and recoveryccs: Software and its engineering Software verification and validationccs: Software and its engineering Software testing and debugging
@@ 1. From opposition to complementarity
  • ‘Seeding Contradiction” (Section 6 and (Huang et al., 2023)), extends the capability to systematic test generation. By deliberately creating failed proofs and running them through Proof2Test, it can produce a collection of tests that satisfy code coverage criteria. It also incorporates loop unrolling to create tests that explore loop behavior and helped provide an assessment of Modified Condition/Decision Coverage (MC/DC), a testing criterion sometimes viewed skeptically but widely used in industry.
@@ 2. The core idea: proofs and their failures
A prover works on the combination of a program and a specification (often expressed as “contracts” as explained in the next section). The starting point for the present work is to take advantage of the technique used internally by many modern provers: they seek one or more “counterexamples” — example inputs falsifying the expected properties. If this attempt fails, the proof succeeds [^1] The duality between failure and success is a general theme of the present work. It follows from the generalization of De Morgan’s laws to predicate calculus: the negation of a universal is the existential of the negation, and conversely. In symbols, $\neg(\forall x\mid p)=\exists x\mid\neg p$. A proof seeks a universal ($\forall$) property; a failed test is an existential ($\exists$) property of the negation.. The present work focuses on the intermediate steps, when the proof fails, meaning the search for a counterexample succeeds; and, more specifically on the counterexamples themselves.
@@
The next application is to use the counterexamples to generate program fixes. Automatic program repair (APR) needs “invariants” characterizing failed executions. Instead of getting them from manually generated tests, the present approach can generate guaranteed invariants; instead of validating fixes through tests, it can guarantee their correctness by running them through the prover. Those are the functions of Proof2Fix (Section 5).
@@ 4. Creating tests from failed proofs
In Fig.  3 (a) the values are small: a two-element array with values 0 and 1. If asked to produce one counterexample, the SMT solver typically yields much larger values. Here it initially produces an array size a.count of 11,800 and values a [11799] ==\, 0 and a [11800] ==\, 5. For disproof purposes (showing that the postcondition does not hold), they are as good as any others. Since the purpose of Proof2Test is to provide concrete evidence helping programmers figure out concretely what is going on, we prefer values that mean something at the human scale. Proof2Test takes advantage of the underlying prover’s ability to generate not just one but several counterexamples (when provided with different seeds). It applies a minimization algorithm (Huang et al., 2022) to generate successive ones until it finds minimal values, such as those of Fig.  3. As with traditional programmer-devised test cases, they enable the programmer to relate to the failure by trying them out, as in Fig. 2, and to see that on loop initialization at lines 10 – 12 Result = 0 and i = 2 so that at line 18 the exit condition evaluates to True, terminating the loop and revealing the fault: early termination prevents the program from getting to the maximum value at position 2. To eliminate the error, it suffices to strengthen the exit condition to permit one more loop iteration: change i $\geq$ a.count to i >> a.count.
@@
Experiment results (Huang et al., 2022) also show that minimization is cost-effective: in most cases, it reduces the values of integer variables by over 80% with an average cost of less than 4 extra reverification runs. Most minimized values are small and relatable: out of 125, 108 are in the range [-2, 2], out of which 58 are zero; others are usually close to some constants appearing in the program.
@@ 7. Related work
The earliest work we know to have applied this idea (Angeletti et al., 2009) generates tests for low-level C programs using Bounded Model Checking (BMC) (Kroening and Tautschnig, 2014), producing test suites with exhaustive branch coverage. A more recent variant, for Java bytecode, is JBMC (Brenguier et al., 2023), where each verification run only activates one assertion at a time, producing one counterexample (the C version (Angeletti et al., 2009) uses compile-time macros, one for each block to avoid actually generating multiple programs); in contrast, the SC framework produces only one program, using a single run-time variable representing the block number. Similarly, (Beyer et al., 2004) extends the software model checker Blast to generate test suites of full coverage with respect to given predicates. BMC-based approaches rely on the correctness of the bound of the execution trace: if the bound is not set correctly, some branches might not be covered, requiring more verification runs to obtain a better bound. The SC approach does not need such mechanisms. An approach (Nilizadeh et al., 2022) that applies ideas of fault injection for generating tests (building on work using counterexamples for program repair (Nilizadeh et al., 2021)) in Hoare-style verification exploits counterexamples produced by the OpenJML (Cok, 2021) verification tool to generate unit tests in JUnit format. These approaches all bear similarities to the present work, which seems to provide the most far-ranging and unifying collection of proof-test combination techniques, from exploiting proof failures to counterexample minimization, test generation covering several code coverage measures and automatic program repair.
Automated test generation is an active research area (e.g. (Gorla et al., 2025)) which in recent years has increasingly used LLM, generally leading to average coverage levels of between 75% to 90%. Proof2Test is much more efficient. LLMs can also serve to generate specifications automatically (Ma et al., 2025), although questions remain: while such specifications might make the code provable, they might also be the wrong specifications, preventing the generation of some relevant tests.
The present work distinguishes itself by the use of Eiffel. While it can be viewed as a limitation, this choice also provides a considerable advantage since it does not require an addition to another language (as JML for Java) but uses an existing (ISO-standardized) language, taking advantage of the built-in contract mechanism and hence directly using the original contract-equipped source code for both tests (with EiffelStudio’s mechanism for evaluating assertions during testing) and proofs. Users of other languages may view the present results as those of “laboratory work”, using ideal conditions so as to open the way for others to extend the results to other environments.